OpenStudy (anonymous):

Verify the identity: cos4x=1-8cos^2x+8cos^4x

OpenStudy (anonymous):

Recall some useful identities: \[\cos2x=\cos^2x-\sin^2x~~\implies~~\cos4x=\cos^22x-\sin^22x\] It'll also help to know that \[\sin^2x=1-\cos^2x\]